### v1

# Designing Robots to Help Women*

Abstract
Robots are being designed to help people in an increasing variety of settings–but seemingly little attention has been given so far to the specific needs of women, who represent roughly half of the world’s population but are highly underrepresented in robotics. Here we used a speculative prototyping approach to explore this expansive design space: First, we identified some potential challenges of interest, including crimes and illnesses that disproportionately affect women, as well as potential opportunities for designers, which were visualized in five sketches. Then, one of the sketched scenarios was further explored by developing a prototype, of a robotic helper drone equipped with computer vision to detect hidden cameras that could be used to spy on women. While object detection introduced some errors, hidden cameras were identified with a reasonable accuracy of 80% (Intersection over Union (IoU) score: 0.40). Our aim is that the identified challenges and opportunities could help spark discussion and inspire designers, toward realizing a safer, more inclusive future through responsible use of technology.
I INTRODUCTION
Within the general area of Feminist Human-Robot Interaction (HRI), the current paper focuses on exploring how robots could be designed to help women to deal with various common challenges, as pictured in Fig. 1.
[…]
Figure 1: Basic concept: robots could be designed to help women related to crime, illness, and day-to-day challenges
As we look toward the future, in which technologies like Artificial Intelligence (AI) and robotics are expected to help people to live better lives, we find similar indications of potential marginalization: For example, reported examples of gender bias in AI include Google preferentially showing high-paying ads to men, Amazon’s hiring system preferring applications from men, Google Translate defaulting to male pronouns, LinkedIn suggesting that female names are mistaken, and Google searches for women often turning up erotic content [2]. As another example, Autonomous Vehicles (AVs) are poised to contribute to safe, efficient transport, but even basic vehicular features such as seat belts, airbags, and crash dummies have up until recently seemingly not taken female body sizes into account, leading to an increased rate of injury in traffic accidents (47% higher between 1998 and 2008 in the US). Furthermore, as Winkle et al. describe, female opinions have also been highly underrepresented in HRI studies, as with Human-Computer Interaction [3]: robot designers are generally not female, and examples of tackling women’s aims and needs appear to be few. (Note: In using the term "women" here, we do not focus on non-binary or gender-fluid cases, although our ideation might also be relevant to some degree for such groups; as well, we follow the typical convention that "sex" refers to biological state and "gender" to social identity.)
One barrier is that the designated design space is so expansive: Women face various challenges that can seem "wicked" and difficult to solve due to their complex, unclear, and interconnected nature. Furthermore, robotic capabilities are ever-expanding; various kinds of robots can be imagined (e.g., drones, wheeled robots, humanoids, exoskeletons, etc.) that could either be carried by a woman, located in an environment, or sent by an authority such as the police or Emergency Medical Services (EMS). Thus, the context suggested the usefulness of following a critical approach that can elicit new insights into both theory and practice.
The remainder of the paper is structured as follows: In Section II we briefly frame and motivate our work in comparison to previous work. Section III describes the speculative prototyping approach we adopted, as well as challenges and opportunities identified. Section IV goes deeper into one specific scenario via a prototype, and Section V discusses results.
In summary, the current paper’s contribution is two-fold:
  • Theoretical. We identify and explore some new opportunities and challenges related to useful tasks for robot designers to help a large user group (women).
  • Practical. We explore one scenario in more depth, reporting on a new proof-of-concept of a robot.
II RELATED WORK
Previous related work seems to have mostly focused on topics outside of the scope of the current paper, such as differences in how women and men perceive robots, or sex robots. For example, women were observed to place more trust than men in the idea of a security robot [4], and sex robots, rather than promoting objectification of women, could actually foster liberalization and sex-positivity [5]).
However, a few studies have explored how robot designers could address challenges specific to women. One work that is highly relevant to the current topic is by Winkle et al., who proposed Feminist HRI, highlighting the importance of examining and challenging power, in considering diverse perspectives, emotions, embodiment, and "low power" users, to achieve responsible design [6]. As well, Winkle and Mulvihill explored how robots could be used to abuse women in a domestic violence (DV) scenario, toward supporting risk mitigation [7]. The authors comment briefly on possibilities of using ambient sensor data or mobile apps to help detect DV, yet advise caution–referring to a study by Cookson et al. that discusses how digital interventions can be overly hyped and under-deliver, have unintended consequences and hidden costs, and alone are no "magical bullet" [8]. A difference with the current work is that these studies did not seek to speculate on what kinds of specific tasks future robots could one day perform to help women, as we seek to do here. For example, helping could also be possible one day for robots in a DV scenario, as perpetrators typically try to isolate their victims and avoid having their crimes recorded or observed–a task that could be complicated by the close presence of a robot outfitted with sensors.
In contrast to the above studies, much work has looked at how robots could use sensing to help people in general. For example, in our own previous work, we explored how a robot could seek to defend a person who is being attacked [9]. We also experimented with combining object detection in RGB and thermal trace detection to detect activities, which inspired our prototype in the current paper [10]. However, these works have not focused on the unique requirements of women. Thus, it seemed like the question of how robots could help women could benefit from some additional attention.
III METHOD
To identify challenges and opportunities, a speculative prototyping approach was followed. Speculative prototyping aims to "drag" potentially important future scenarios from the foggy realm of imagination closer to the real world, visualizing in a thought-provoking, clear, visceral manner, as a means to begin a process of critical exploration [11]. Thus, rapid ideation sessions were first conducted to identify challenges affecting women. Next, opportunities were identified. Ideas were recorded without judgement, then afterwards merged and reworked with priority attributed to those that seemed valuable, different, and requiring further study. (In doing so, we aimed to roughly align ourselves with Suvin’s concept of "cognitive estrangement", described by Bartolotta in terms of seeking to capture "fictional world-affordances imbued with cognitive potential" [12].) Thus, some representative scenarios were chosen and transformed into sketches. Furthermore, one of these sketches was further developed into a physical prototype, which was used to also gain some insight into practical concerns.
III-A Challenges Faced by Women
What then are some of the challenges that seem to be predominantly faced by women? In considering this question here, we adopt a zemiological perspective; i.e., a broad view that can encompass various potential harms, in relation to health, crime, and day-to-day activities.
III-A1 Crimes
Crimes that disproportionately affect women include voyeurism, sexting without permission, stalking, indecent exposure, domestic abuse (including intimate partner abuse), rape (including date rape, acquaintance rape, marital rape, gang rape, and war rape), molestation (grabbing, groping), child abuse and grooming, forced prostitution (including sex trafficking, sexual slavery, forced marriage, revenge porn, and forced participation in the production of pornography), and other forms of sexual assault, harassment, bullying, and violence [13]. [^4] https://pcar.org/about-sexual-violence/adults We note that these terms are not exclusive but often contain some overlap. As well, crimes can be further detailed in terms of individual actions. For example, physical violence can involve slapping, pushing, hitting, kicking, hair-pulling, choking, burning, or assault with objects such as knives or guns. Also, even such non-sexual violence can relate to sex, since women who are smaller can be seen as easier targets [14].
Such distressing experiences are not limited to any one country. For example, a report from 2002 indicated that two thirds of Swedish women have been sexually harassed or experienced violence [15]: In more detail, over half of women have been sexually harassed, and around half have experienced violence (a quarter physical violence, one in three sexual violence), including around a quarter of young women 18-24 just in the past year; one out of every five women has also been threatened by a man. Numbers are also disturbing in the US, where around two thirds of women have experienced violence, and around one fifth of women have been raped, usually by someone they knew [13]. In some regions, such as West or South Asia and Africa, women can also face risks of female genital mutilation, honor killings, stonings, acid throwing, blinding, persecution due to alleged witchcraft, and forced suicide (e.g. Sati, in which a widow is forced to die on her husband’s funeral pyre [^5] https://en.wikipedia.org/wiki/Sati_(practice)).
The crimes above can be categorized as occurring remotely, or in proximity. Of the first non-contact category, voyeurism, also related to scopophilia, upskirting and peeping, could be one useful problem for initial exploration: In contrast to more serious crimes where the "stakes" can be extremely high if a mistake is made, voyeurism has been called a "harmless wrong", as victims often don’t know they have been victimized and thus do not suffer trauma [16]. Moreover, current legislature regarding filming without consent seems to support technological approaches to limit video voyeurism, and it is also arguably clear how voyeurism can be stopped, by preventing a criminal from viewing and filming victims. As well, voyeurism is an important, widespread problem: Although actual numbers are difficult to predict, voyeurism has been described as an epidemic in Korea, with over 6000 cases reported each year [17]. When footage is spread, the result can be "social death", causing some women to attempt suicide. For such reasons, in 2018, approximately 20,000 women marched to protest spy cameras, with 200,000 signing a petition, which resulted in a government plan to hire 8000 workers to tackle the problem. As well, in Sweden, in an anonymous survey of 2,450 randomly selected 18–60 year-olds conducted in 1996, 7.7% of respondents described deriving sexual satisfaction from spying on others having sex [18]. Thus, it could be useful for women if voyeurism could be more easily detected and prevented.
Of crimes that occur in proximity, two kinds that seem more difficult to tackle, but highly important, include DV and rape. For example, in Sweden, there are typically no witnesses to DV, as the attacks occur "behind closed doors" [19]. As such, various statistics on DV can be found, but the real numbers, and hence the true extent of this problem, are unknown. As well, hospital records are often generated as a result of attacks, but police are usually not allowed to access these data; victims can also feel retraumatized and uncared for when interacting with healthcare staff.
Furthermore, recent discourse has highlighted difficulties in Sweden related to immigration, integration, and crime [20]: Especially victims coming from other cultures can find themselves in a position of weakness, segregated, and like they have no one to talk to: They might not know the local language (Swedish), be allowed to go outside, or be aware of Swedish norms (i.e., no one might explicitly tell them that is not okay for a man to beat a woman). They might also visit religious centers that reinforce non-Swedish views, like that they must obey the men in their families, or that they could be punished or killed for having a local boyfriend. For example, views in Africa on DV can be startlingly different, with acceptance of wife beating at 77% in Mali and Uganda; overall, "51% of African women report that being beaten by their husbands is justified if they either go out without permission, neglect the children, argue back, refuse to have sex, or burn the food." [^6] https://blogs.worldbank.org/en/africacan/domestic-violence-and-poverty-in-africa-when-the-husbands-beating-stick-is-like-butter
Another key problem with DV is that repeat offenses are common [21]. For example, officers might pick up a criminal on Friday for beating a woman, then release him on Saturday, in a pattern that repeats itself each week. The perpetrator as well can escalate, beating harder and harder each time, which can end in murder. Some mechanism is needed to break such damaging spirals.
[…]
Then, what might be needed? As above, one challenge is getting information to the victims, so they are no longer alone. Also, women who are victims need to be protected; something should be done for them without them having to start the process.
Another important problem to combat is rape, which shares many similarities with DV: Rapes are usually conducted by perpetrators known to the victim; the crime is committed in close contact; its repercussions are among the most serious of crimes (leading to chronic health problems such as PTSD, MDD, and substance abuse); help-seeking victims often experience a "second rape" in being repeatedly interrogated, doubted, blamed, warned, and discouraged; and a chronic failure to report or investigate rape has been noted [24] [25]. In some countries, like the US, clearance rates for rape also appear to be low (e.g. in the 60 percent range, compared to Japan in the high 90s, with rapes being about 100 times more common) [26].
Currently, in Sweden, one obstacle to preventing rape through adopting new technologies comes from the legal system. For example, the European Parliament, in its recent "Artificial Intelligence Act", has banned real-time detection, on the basis of the potential for mistakes. [^11] https://www.europarl.europa.eu/news/en/press-room/20240308IPR19015/artificial-intelligence-act-meps-adopt-landmark-law More concretely, in an urban environment with surveillance cameras, it is not permitted to detect rape in real time and alert police. Rape must happen first–everyone is required to wait–then the images can be used, after judicial approval is obtained. As such, currently much of the potential use of AI is lost, and the integrity of victims is not upheld. Thus, here the challenge seems to be not merely technical but also political. While it’s clear that current recognition systems are imperfect, various examples of working real-time systems exist, not just in pacemakers and brakes, but also employing computer vision, from Tesla cars doing real-time inference from cameras, to real-time pixelization of faces on television. Furthermore, a basis exists for how to handle imperfect recognition in standards such as SOTIF (ISO/PAS 21448). [^12] https://www.iso.org/standard/77490.html Thus, an example of a prototype of how robots could detect and prevent crime in real-time at an early stage could also be useful as a way to affect thought regarding legislature.
Another problem involves bottlenecks in processing evidence. Disturbingly, it seems there are hundreds of thousands of untested "rape kits" in the US, some of which are not processed even after 30 years [27]. [^13] https://usafacts.org/articles/how-many-rape-kits-are-awaiting-testing-in-the-us-see-the-data-by-state/ This forensic evidence, which is costly and requires hours of invasive handling of victims, is sometimes not sent onward for analysis or cannot be handled due to insufficient resources. Various potential causes have been put forward such as budget cuts; victim-blaming and bias against women and victims of sex crimes; slow workflows due to fear of contamination and inefficient or redundant testing/auditing requirements; lack of a tracking system for forensic evidence; and misunderstandings of processes surrounding rape kits. In some cases, rape kits have been destroyed before testing without notifying victims. In some other countries, lack of rape kits or trained examiners can also be a problem. The result can include decreased community trust in police, as well as missed opportunities to identify serial perpetrators and prevent new rapes and other crimes–as rapists are often also guilty of other crimes such as burglary, assault, and murder.
III-A2 Health Challenges
Health problems that disproportionately affect women include breast cancer, gynecological/period-related problems (e.g., premenstrual syndrome (PMS)/premenstrual dysmorphic disorder (PMDD), perimenopause/menopause, dysmenorrhea, endometriosis, ovarian and cervical cancer), complications due to pregnancy (including difficult childbirth and abortion, as well as perinatal depression), eating disorders, anxiety, migraines, osteoporosis and autoimmune diseases [28]. [^14] https://online.regiscollege.edu/online-masters-degrees/online-master-science-nursing/womens-health-nurse-practitioner/resources/health-issues-specific-womens-health/
In general, women in the European Union (EU) report poorer health and poorer mental well-being than men, and are more likely to have health limitations over their lifetime [28]. This might be why women are also more inclined than men to attempt suicide [29]. It also often takes women longer to be diagnosed with illnesses, possibly due to gender bias. [^15] https://www.weforum.org/agenda/2024/02/womens-health-gap-healthcare/
III-A3 Other Day-to-Day Challenges
The UN lists eleven largest hurdles for women’s equality by 2030, which includes lack of women in leadership positions, workplace discrimination, and an imbalance in unpaid care work [30]. For example, the report mentions typical rates of 20-30% of managers being female, women earning globally 51 cents per dollar earned by men, and an estimated extra two hours of unpaid care work each day that women will still be expected to spend by 2050.
Physically, women are also on average smaller, shorter and weaker than men (possessing less muscle mass and bone density and more body fat) [31]. This might not merely affect various day-to-day tasks such as reaching high-up items, carrying heavy objects, and opening tight containers; for example, height also affects perceived dominance, which might explain why most Chief Executive Officers (CEOs) of Fortune 500 companies tend to be tall men [32]. Another difference is often seen in clothing, which has repercussions related to health and crime. Some clothes can be used by criminals as an excuse to commit crimes, such as purses that seem easy to snatch, or revealing clothing such as skirts, which can be seen to invite upskirt photography or rape [33]. In Sweden, perpetrators sometimes also come from countries with different cultures, where they might never have seen any of their female relatives (e.g., mother or sister) naked, or women in light clothes; they might also have never been explicitly told by anyone how things are in Sweden.
As well, certain clothes can be impractical, such as high heels, which can lead to foot pain or tripping, or even merely a lack of clothes with (large) pockets, leading to increased use of handbags, which can be stolen [34]. In some countries, women can be hassled or killed if they do not clothe themselves in a certain way (e.g., by Mutawa "religious police" or relatives, if a hijab is deemed to not be sufficiently set); 15 girls were reportedly burned alive in 2002 in an incident when the Mutawa beat girls trying to escape from a burning school without their hijabs [35]. Women are also more targeted by the fashion and cosmetics industries, which are among the biggest international businesses, and encouraged to spend more time on physical appearance and beauty products (e.g., seeking to avoid tanning in Asia) [36]. As well, products are sometimes sold at higher prices targeting women that only differ in packaging, fragrance, or color (e.g., razors similar to those sold to men, but just in pink color); such gender-based price segmentation has been described as a "pink tax", and has also been observed to some extent in Sweden [37].
III-B How Robots Could Help Women
III-B1 Robots Helping to Fight Crime
In regard to voyeurism, hidden spy cameras could be detected. For example, Sami et al. describe a smartphone app that uses laser/time of flight to detect hidden cameras [38]. Yu et al. used a thermal camera to detect hidden cameras via heat [39]. Here we also build a prototype drone aimed at dissuading voyeurism by detecting hidden cameras. One difference with the concept for our prototype is that both approaches above in their current form still require a human to think of installing and running an app and to move around an apartment searching for devices themselves, which might not be easy in hard-to-reach or high locations. Another difference is that our proposed approach also involves object detection.
In regard to proximate crimes like DV, rape, molestation, and stalking, robots could seek to prevent woman from being targeted and perpetrators from getting away (i.e., preventing the perpetrator’s ability to carry out an attack and ensuring justice is served). Deterrence could be partially achieved by reducing times in which young women must be alone, e.g., via the company of a robot that can record evidence and call for help. Robots could accompany women if there is a perceived risk, such as in a home where DV has occurred, or if she needs to move through a dangerous part of a city, like a parking lot at night. Robots like drones already could seem threatening, taking off, buzzing loudly, flying erratically, and threatening collision, if someone gets too close. This could be like suddenly summoning a loud barking dog to one’s aid, that can otherwise be turned off and carried easily, and could also help to avoid purse-snatching or upskirting. A robot could also indicate a woman’s personal space by projecting a circle of light around her, or shine a light on a stalker following a woman. One related example in the literature exists, of a "spider dress" designed by Anouk Wipprecht that inflates, extending mechanical arms, by monitoring proximity and a woman’s breathing; another variant releases smoke. [^16] https://medium.com/@intel/is-that-a-spider-on-your-dress-or-are-you-happy-to-see-me-da25075314b9 A robot could also prevent criminals from getting too close not only to a woman, but also to objects; e.g., it could prevent someone from getting close to a woman’s drink to insert a sedative. If a woman in a man’s presence suddenly seems incapacitated, this could also indicate that the robot should call for help or separate them. Intent to attack (e.g., hands balled into fists or hidden, getting close, angry/loud language, etc.) could also be detected.
In a more complex scenario, a remote robot might be alerted (e.g., by detecting nearby victims’ screams [40]), or be sent by authorities. In such a case, advanced capabilities might be required to distinguish true attacks from jokes, skits, or play-fighting, and make judgements about "distinction" and "proportionality": A robot might need to identify which person is the victim and which is the perpetrator–a highly challenging task when multiple people are present. As well, decisions about the level of force to use might include some analysis of force differentials, backdrops and crossfire. Additionally, age detection could also be used to detect child abuse or forced marriage, and flying robots or soft robots could be useful to enter closed buildings in which trafficking might be occurring.
During a crime, the robot could target the assailant from a difficult angle (e.g., from behind or from above for a drone) with pepper spray or laser to dazzle the assailant’s eyes, while emitting loud sounds and bright lights to seek help. As well, the robot could try to mark the attacker or their vehicle with paint, and record license plates, or interfere with their movement. The robot could also provide advice during the attack: e.g., to drop to the ground to be harder to move and to avoid being taken to a "second location".
Robots could also help post-hoc, if an attack could not be prevented. For DV, an AI tool could track hospital records to assess risk, if current laws are changed. Federated learning, e.g., downloading updates to AI models for detecting DV without uploading sensitive or restricted data, could be one way to ensure that private data are not misused. After a rape, a robot with a sterile compartment could somehow try to facilitate rape testing immediately, avoiding waiting time at hospitals, also since robots lack the DNA of a human investigator (there might be less risk of contamination). As part of this, robots could also visually analyze victim’s bodies. Fernandes et al. proposed and evaluated a deep learning approach to identify genital lesions that could be evidence of rape, using a dataset of roughly 400 images collected by the Southern Denmark Sexual Assault Referral Centre (78 from non-consensual and 316 from consensual intercourse) [41]. Robots could also help to search for a hostage or corpse, or track fleeing attackers. In the undesired case that a woman has been murdered, robots are being built to do forensics on corpses, like the Virtobot system [42].
III-B2 Robots Helping to Support Health
To detect health problems, a robot could generate three-dimensional (3D) scans of a person using a Lidar device. For example, scans can already be done using Lidar devices in IPhone Pros or IPads. The benefits are:
  • Current methods sometimes haven’t been updated for a long time and could be upgraded. For example, tape measures are used to measure belly girth for pregnancy, which can be inaccurate.
  • People can measure themselves outside of infrequent doctor visits and have continuous control over their own health data (relating to data democratization and data-driven care). For example, in some parts of Sweden, breasts x-rays can be done every five years, and in between, patients are expected to wait.
In more detail, such 3D scans could be used to detect breast cancer, anorexia, and abnormalities:
  • A change in a breast (e.g. due to a lump) could indicate potential cancer.
  • Weight loss in a short time could indicate anorexia.
  • A scan could help someone to understand that it is normal for women to not all be shaped like photo models or porn stars (i.e., there is much variance). For example, this is a goal in the initiative of Visual Sweden called "Visual Vulva". [^17] https://www.visualsweden.se/en/aktuella-projekt/ar-jag-normal/
Where simply scanning is insufficient, robots could intervene in a more complex manner. For example, robots have been designed to carry out clinical breast examinations. [^18] https://www.bristol.ac.uk/news/2023/october/new-robot-could-help-diagnose-breast-cancer-early.html As well, AI systems have been designed for diagnosing ovarian cancer [43], which can be tackled via robotic surgery [44]. Anxiety and depression (which are sometimes exacerbated by loneliness) could be aided by cuddly companion robots such as Paro [45]. To help women with their periods, a robot could apply heat or massage to minimize cramping, and remind about drinking or offer water to minimize bloating. To further help women with eating disorders, a robot could try to prevent binge eating (e.g., hide food or discourage purchases), provide dietary advice, reinforce a positive body image, stop excessive exercise by positive distraction, or cook with a person.
To help women during pregnancy, a robot could perform duties like a midwife, checking baby position (upside down or not) and heart rate, etc. For women experiencing morning sickness, a robot could detect patterns to identify causes of ill feeling (e.g., if certain foods are a problem) and make healthy food. Various robots capable of making food exist, including a prototype we developed that checked how salty or sweet its cooking was to ensure healthiness. [^19] https://www.youtube.com/watch?v=6eqmtkOm-Tk As well, a robot could detect vomiting, and clean when needed. A robot could also detect fainting and either catch the woman or call for help if needed. The robot could also find places for a woman to rest when walking, like benches in shadow, and fan or cool her.
III-B3 Robots Helping with Other Day-to-Day Challenges
Regarding other day-to-day challenges, exoskeletons or other tools could help to level the playing field, by allowing anyone to be large, tall, and strong. For example, this might allow smaller women to become accepted as leaders, or to be accepted as equals, receiving equal pay, in physically intensive jobs such as construction or policing. As well, smaller women could reach high-up items, carry heavy loads, and open tight packages–and maybe even better defend against attackers, helping them to fight back and be too heavy to knock down or abduct. Furthermore, alloparenting robots capable of raising children in a good way could also help [46]; for example, when pay is unequal due to the fear that women will leave work to take care of children, or women are overwhelmed with unpaid care.
III-B4 Sketches
From the pool of scenarios, five were selected and transformed into sketches, as shown in Fig. 2.
Figure 2: Sketches: (a) Nazar - A helpful flying "eye" could sense "evil" such as hidden cameras in a woman’s room, (b) Hero - A robotic barrier could seek to ensure that people intending harm cannot enter a woman’s personal space, (c) Midwife - A helper robot can take care of pregnant women, (d) Allo - A robotic baby carriage can try to help mothers by taking care of repetitive child raising tasks, (e) Equa-skeleton - Exoskeletons could create equal conditions
IV PROTOTYPE
Theoretical ideas alone sometimes miss practical realities that can be exposed by prototyping. To gain further insight into one of the scenarios, proposed in the first sketch, a proof-of-concept was created, as shown in Fig. 3, in three steps: Preparation. First, a mock-up environment was created containing some objects with hidden cameras and distractor objects intended to make the detection task more difficult. In total, 22 objects were gathered, a similar number to previous papers: Toilet Paper Roll, Book, PET Bottle, Pouch (Snowman), Toy Car, Toy Mammoth, Pill Bottle, Toy Frog, Wine Box, Lamp, Toy Box, Clock, Toy Horse, Gum Bottle, Toy Egg, Glasses, Banana, Sponge, Sunscreen, Toy Buzzer, Medicine Box, Paper Roll. Some standard Logitech/Plexgear webcams were hidden within five objects: Toilet Paper Roll, Toy Car, Pill Bottle, Wine Box, Lamp. Furthermore, a simple OpenCV program showing the video feed was run for each camera to simulate recording. [^20] https://opencv.org/
Robot Motion. Second, we recorded some footage of the drone, a Ryze Tello, flying in front of the objects. We explored various forms of control, e.g., manually controlled the drone via the Ryze Tello smartphone app, as well as using the EasyTello python library [^21] https://pypi.org/project/easytello/ to issue commands and obtain video, and detecting ArUco markers that could be used to guide the robot or for distance estimation.
Hidden camera detection. Third, we explored how spy cameras could be detected using a thermal camera/RGB camera. Since the cameras weighed too much (90g without battery, vs. 80g payload for the drone), for our initial exploration, data were obtained with the camera placed on a desk approximately 1.5m in front of the mock-up environment.
Fig. 4 shows this basic process: Raw RGB and thermal data were obtained using, respectively, a Sony IMX219 8-megapixel sensor, and an inexpensive 80 x 60 FLIR camera capable of detecting heat in the range of 8–14 µm. After calculating a mapping, the thermal camera was used to find warm areas in view that might arise from a hidden camera. Simplified thresholding was conducted to derive a mask. Next, an algorithm was run to detect which objects might be responsible for the warm spots. The aim was to reduce the area that needs to be searched and make it easier to find the cameras. YOLO version 8s from Ultralytics was run on the RGB image with the confidence parameter set to 0.1 to detect the locations of objects as a set of bounding boxes. [^22] https://www.ultralytics.com/yolo YOLO uses deep learning (a convolutional neural network) to detect objects, where the confidence threshold handles non-maximum suppression. Finally, bounding boxes enclosing heat traces were selected as the output of the system, representing objects potentially concealing cameras, that a robot could either remove or show to a human.
As a result, 5/5 of the cameras’ heat signatures were clearly visible after thermal thresholding, resulting in five detected contours. Next, boundary boxes were detected for 4/5 (80%) of the objects enclosing hidden cameras (although overall, only half, or 11, or the 22 objects were detected by YOLO). Finally, the agreement between the ground truth and system output regarding locations of objects concealing hidden cameras was calculated. The average Intersection over Union (IoU) was 0.401 (min=0.0 for the Toy Car which was not detected as an object, max=0.881 for the Wine Box).
Figure 3: Prototype concept: a drone could check a room for hidden cameras
Figure 4: Image Processing: (a-b) raw RGB and thermal images, (c-d) intermediate output from YOLO and threshold on thermal traces, (e-f) comparison of overall groundtruth–where purple boxes with red asterisks are the target, and distractor objects are labelled in other colors–and the system output, with groundtruth in blue, output in green
V DISCUSSION
In the current paper, three main categories of challenges that disproportionately affect women were identified: criminal, medical, and day-to-day. From these challenges were born ideas for opportunities, of how designers could use robots to better women’s lives. Five scenarios were visualized via sketches. Furthermore, the first sketch was implemented as a prototype, of a drone system that uses a thermal/RGB camera to detect hidden cameras. A summary of this work is also available via an online video. [^23] https://youtu.be/7slpfGD1sEU
Regarding the simplified prototype intended to detect hidden cameras, the result of 80% accuracy (IoU 0.401) seemed reasonable, given the challenging mock-up and the simplified approach. As a result of some simplified experimentation, we observed that lowering the confidence parameter allows YOLO to detect more bounding boxes, such that all five objects of interest are detected; a demerit is that reducing the confidence threshold results in many more objects detected, which can result in overlapping bounding boxes and increased complexity. Also, although the IoU score is imperfect, we believe this should not be a deal-breaker since a robot or person is not limited to checking only the exact inside of each bounding box but can also check the vicinity, as long as they have a rough estimate of where the camera might be.
Also, YOLO can be used not only to obtain bounding boxes, but also to recognize objects. This could be useful in various ways. For example, some natural heat sources like ovens or people could be removed from consideration. For others, such as lamps or computers, a model of where such devices normally become warm could be used to detect anomalies caused by hidden cameras. As well, a label could be directly provided to the human in regard to which objects are expected to contain hidden cameras. However, in our simplified exploration, object recognition appeared to have been challenging, possibly due to illumination, the cluttered mock-up environment, or camera limitations. As can be seen in Fig. 4, four objects were recognized correctly, two teddy bears, a clock, and a book, whereas seven objects were recognized incorrectly, as a sink, cup, bottle, donut, cellphone, knife, and book. In such a case, a robot might be able to avoid having to recognize objects by instead showing people where cameras are expected to be located on a screen (e.g., either a screen on the robot, or by sending a picture to a person’s smartphone).
Along the side, one thing we observed during preparation was that hooking up all five cameras to an old desktop resulted in crashes, possibly due to overwhelming the bandwidth on the same USB bus, so for subsequent attempts several computers were used. Another observation was that our drone, like other typical drones used by the community such as Crazyflie, would not able to detect objects on the ceiling due to its camera placements, as it only has two cameras, that look forward and below respectively.
V-A Limitations and Future Work
The current work is limited by various factors, including a bias toward focusing on Sweden, given our location. As well, our ideas come only from us, whereas future studies should gather ideas from a larger pool of female participants, in different countries. Our ideas are also speculative, and thus by intention largely ignore some current realities: e.g., that robots are costly, difficult to maintain and fix, and limited in battery life and capabilities. For the prototype, results are limited by the controlled lab environment (e.g., detecting of expensive, upper range spy cameras from blurry images of a moving drone might require more complex methods).
Furthermore, new technologies create not only opportunities, but also potentially new crimes. Robot designers should consider ahead of time how robots could be misused, and factor in such questions. For example, how to ensure that times or images recorded by a robot are accurate. Another question is if robots could inadvertently harm the women that they seek to help. For example, wind from a drone’s propellers could disturb a crime scene, stirring up dust, erasing evidence or introducing contamination. Or, robots could make it harder for women to escape or hide (e.g. if a costly or large robot crashes, blocks an escape route, or betrays a woman’s location). As well, the concepts in the developed sketches could also be potentially misused to harm women. For example, the camera detecting prototype could be used by criminals to ensure that their cameras are well-hidden and will not be easily detected. Or, a robot that guards a woman’s personal space could be hacked to slip a sedative into her drink. Furthermore, (although it is estimated that false accusations, e.g., of rape, tend to be not so common) criminals could also potentially misuse robots to get others in trouble (e.g., getting an accomplice wearing a mask to trick the robot into thinking an innocent person committed a crime). Thus, one important area of future work will involve refining such scenarios, such that for example robots could be used to facilitate legal processes by delivering new or better forms of evidence in courts.
Another question regards who will control women’s ability to protect themselves with robots. Although governments introduce various rules and regulations, they might not "set the pace", since criminals can ignore rules, and people often use apps, which have their own terms. While users cannot be aware of all sets of terms for the apps they use, even if rules of technology providers might be in conflict, it seems like women themselves could set up the apps and robots they need for protection, e.g., detecting rapes in real-time with a drone. And, if such usage becomes widespread, it could motivate modifying existing restrictions. Therefore, we believe that prototyping solutions in this area could be a useful showcase for what AI can do to help people.
In summary, we believe that further exploration of the challenges and opportunities that arise in designing robots to help women could help us to realize a safer, more inclusive future for women, and thereby for all of society.